OpenStudy (anonymous):

Our cost is $55. We want to resell the product so that we make a 30% profit. Find the resale price.

OpenStudy (anonymous):

Why is the answer not simply \[55(1+.30)=price\]

OpenStudy (anonymous):

Why is the solution really...\[\large \frac{55}{100-30 \over 100}=price\]

OpenStudy (anonymous):

im not sure...i thought you would find 30% of 55 and then add that to $55 to get the resale price now im not sure...

OpenStudy (anonymous):

Your solution is \[55(.30)+55=55(1+.30)\] but thanks anyway

OpenStudy (anonymous):

ok sorry

OpenStudy (anonymous):

I never took finance so I'm trying to wrap my mind around this

OpenStudy (anonymous):

oh well neither did i haha sorry im in 10th grade

OpenStudy (anonymous):

It's based on how they calculate the percentage. They want\[\frac{price-55}{price}=.3\]I don't know why that's the way they do it, but just go with it I suppose.

OpenStudy (anonymous):

That actually makes a lot of sense. Thank you

OpenStudy (anonymous):

\[\frac{p-55}{p}=.3\]\[p-55=.30p\]\[55=p-.30p\]\[55=p(1-.30)\]\[p=\frac{55}{.70}\]\[:)\]